find the current in a circuit if the power is 500 w(watts) and the resistance is 25 ohms. round off your answer two decimal place. use formula i=p/r
Answer by Alan3354(69443)

find the current in a circuit if the power is 500 w(watts) and the resistance is 25 ohms. round off your answer two decimal place. use formula i=p/r
------------
Using i = p/r
i = 500/25 = 20 amperes
Any reason you couldn't do that?
--------
Actually P = i^2r, not ir
--> i = sqrt(20) =~ 4.472 amps